GAS M&O meeting 28 October 2019
Present: Beatrice, Roberto, Michal, Fred, Patrick, Kacper, Andrea, LPh, Mara, Gianluca, Sabino, Federico, Fabio
Indico: https://indico.cern.ch/event/859067/
ALICE #no elog
Reminder:
-
- Dummy chamber: Arturo will install the pipes in January, gas team will do weld
- exhaust pressure regulation does not work
- mixer pressure regulation does not work: Zimmerli not installed
- TPC
Reminder:
-
- LPh to prepare a summary on how the system is now —> this week
- TRD
- Rack 66 in error and not selected —> LPh to check
Reminder
-
- Purifier: to review if keep the safety valve or the rupture disk
- Purifier: safety valve at the input where we have Ar or Ar/H2 at high pressure —> LPh to check with safety
ATLAS #2036
Reminder:
-
- After the power-cut there was not communication with PLC: damage of differential circuit breaker —> to repair
- MDT
Reminder:
-
- Rack 68 (ch 17): one channel is keeping going to zero —> try to change the ELMB
- General
- Power-cut due to smoke alarm in SGX1 due to external work
- RPC
- Andrei is finishing the pipes
- This week to send everything to cleaning
CMS #2034
- RPC
- OFF
- Anton asked to prepare all dummy chambers
- Rack 85: problem of pressure sensors in error: problem in the cable (UX-US) —> contacted Anton to change it
- Feedback valve in exhaust to check —> Patrick: the N2 was closed
- Commissioning of the software on-going
- To add filter for pump
- DT
Reminder:
-
- Test for power cut and re-start for the exhaust line —> to do other tests
- CSC
Reminder:
-
- Warning to set-up —> Beatrice/Michal/Roberto
- GEM
- Commissioning started: elog 2017
- Start planned for 4th November: connection of 12 channels GE1/1- from tomorrow
Reminder:
-
- To modify the réseaux socket for the GEMPLC (now it is with TOTEM) —> Andrea
- Control rack: to modify power supply and electrical cable (still with TOTEM) —> Patrick/Michal
- 1 chamber in purge in CO2
- R134a recuperation
- Added line of Air for the R134a recuperation
- Add line for iC4H10 exhaust this week —> on-going
- To check for supply of 220 V + 400 V (Lauda) —> Patrick
- To understand logic in case of presence of leak/iC4H10 —> Roberto/Nicola Broca
LHCb #2035
- RICH1
- RICH2
- need to instal filter but it will arrive mid November: to look for another place before the filter of the exhaust
- GC
- PC is working but not the GC —> today company will come to check
- SciFi
- Start electrical work
- Need to stop all system to modify N2 line —> this week
PIQUET
- elog 886
- Alarm for HH temperature in cloud
- elog 885
- ALICE TRD too long injection time: Andrea changed the mixer flow injection to 5h but he said that alarm was still active
- LPh to verify why we inject for such a long time
GENERAL
- Discussion about to change filter in exhaust with 1 um everywhere
- R134a presentation by Gianluca
- Presentation of 3D drawings of CPV Humidifier by Sabino
- LPh confirmed that all the material is available at the workshop except one festo volume
- Presentation about regulation valves for CMS RPC by Sabino
There are minutes attached to this event.
Show them.